OpenStudy (anonymous):

Solve by completing the square: x^2-8x+2=0

OpenStudy (anonymous):

\[x^2-8x=-2\] is a start then what is half of \(8\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok then we go to \[(x-4)^2=-2+4^2\]

OpenStudy (anonymous):

what is \(-2+4^2\) ?

OpenStudy (anonymous):

16.2

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[-2+4^2=-2+16=14\]

OpenStudy (anonymous):

now we are at \[(x-4)^2=14\] take the square root, don't forget the \(\pm\) and get \[x-4=\pm\sqrt{14}\] then add \(4\) to both sides and finish with \[x=4\pm\sqrt{14}\]

OpenStudy (anonymous):

ok thanks. Can you help me with few more? @satellite73

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Solve by completing the square: x^2 +6x-3=0

OpenStudy (anonymous):

same idea as last time start with \[x^2+6x=3\] then what is half of \(6\) ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

k we go right to \[(x+3)^2=3+3^2\] now (careful) what is \(3+3^2\) ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

got it so \[(x+3)^2=12\\ x+3=\pm\sqrt{12}\\ x=-3\pm\sqrt{12}\]

OpenStudy (anonymous):

ok thanks. and this one too. 2x^2-5x+6=0

OpenStudy (anonymous):

this is not so fun to complete the square with easier to use the quadratic formula do you know it?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok then for \[2x^2-5x+6=0\] use the quadratic formula with \[a=2,b=-5,c=6\] to get your answer

OpenStudy (anonymous):

ok ill finish it. can u help me with few more? I would really appreciate it

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

k

OpenStudy (anonymous):

(5x-1)^2 =24

OpenStudy (anonymous):

we are half way home at the start begin with \[5x-1=\pm\sqrt{24}\]

OpenStudy (anonymous):

then add one get \[5x=1\pm\sqrt{24}\] and finally divide by \(5\) to get \[x=\frac{1\pm\sqrt{24}}{5}\]

OpenStudy (anonymous):

you can also write this as \[x=\frac{1\pm2\sqrt6}{5}\] if you like

OpenStudy (anonymous):

thanks!!! 3x^2 +2x=-4

OpenStudy (anonymous):

start with \[3x^2+2x+4=0\] and use the quadratic formula with \[a=3,b=2,c=4\]

OpenStudy (anonymous):

and next? plz

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you mean how to you use the formula?

OpenStudy (anonymous):

solve everything with the answer

OpenStudy (anonymous):

\[3x^2+2x+4=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=3,b=2,c=4\] you get ' \[x=\frac{-2\pm\sqrt{2^2-4\times 3\times 4}}{2\times 3}\]

OpenStudy (anonymous):

since the number inside the radical is negative, there is not real solution for this one

OpenStudy (anonymous):

so what will be the anwer

OpenStudy (anonymous):

no real solution?

OpenStudy (anonymous):

@satellite73